# Two point functions and quantum fields in the anti-de Sitter universe [^1] Dedicated to the memory of Henri Epstein

### v1

@@ 7.1 Plane waves again
We are now in position to give a global definition of the anti-de Sitter plane waves. Given any complex number $\lambda$ the natural domains of definition for plane waves in the AdS universe are ${\cal Z}_{+}\times\overline{{\cal C}}_{-}$ and ${\cal Z}_{-}\times\overline{{\cal C}}_{+}$:
@@
A previous attempt [18] to introduce global waves was based on a local $i\epsilon$-prescription. Here, in their global domains of definition, the waves are univalued, holomorphic solutions of the Klein-Gordon equation with complex squared mass $m^{2}_{\lambda}=\lambda(\lambda+d-1).$
@@ 8 Klein-Gordon fields
Klein-Gordon fields display the simplest example of the analytic structure explained in Sect. 5. Their two-point functions are distributions on $\widetilde{}AdS_{d}\times\widetilde{}AdS_{d}$ satisfying the AdS Klein-Gordon equation w.r.t. both variables
$(\square+m^{2}_{\lambda}){\cal W}(x_{1},x_{2})=0,$ (8.71)
where the complex squared mass is parametrized by a complex parameter $\lambda$ as in Eq. (6.50). For any such $\lambda$ there are two invariant maximally analytic solutions of the above equation, satisfying all the properties specified in Section 5, apart from the positive definiteness which may hold or fail:
$W^{(d)}_{\lambda}(z_{1},z_{2})$ $=$ $W^{(d)}_{\lambda}(z_{1}\cdot z_{2})={e^{-i\pi{d-2\over 2}}\over(2\pi)^{d\over 2}}((z_{1}\cdot z_{2})^{2}-1)^{-\frac{d-2}{4}}Q_{{d-2\over 2}+\lambda}^{d-2\over 2}(z_{1}\cdot z_{2})$ (8.72)
$=$ $\frac{(z_{1}\cdot z_{2})^{1-d-\lambda}\Gamma(d-1+\lambda)\,_{2}F_{1}\left(\frac{d-1+\lambda}{2},\frac{d+\lambda}{2};\frac{d+1+2\lambda}{2});{(z_{1}\cdot z_{2})^{-2}}\right)}{2^{d+\lambda}\pi^{\frac{d-1}{2}}\Gamma\left(\frac{d+1}{2}+\lambda\right)}$ (8.73)
$W^{(d)}_{1-d-\lambda}(z_{1},z_{2})$ $=$ $W^{(d)}_{1-d-\lambda}(z_{1}\cdot z_{2})={e^{-i\pi{d-2\over 2}}\over(2\pi)^{d\over 2}}((z_{1}\cdot z_{2})^{2}-1)^{-\frac{d-2}{4}}Q_{-{d\over 2}-\lambda}^{d-2\over 2}(z_{1}\cdot z_{2}).$ (8.76)
@@
In the above formulae Q{Q} denotes the associated Legendre function of the second kind; it is a meromorphic function of the indices [19]; as function of the complex variable $z=z_{1}\cdot z_{2}$, it is analytic in the cut-plane
$\Delta_{1}=\{{\bf C}\setminus[-\infty,1]\}.$ (8.77)
As explained in Section 4, the scalar product can be lifted to a map of $\widetilde{\cal Z}_{-}\times\widetilde{\cal Z}_{+}$ onto the covering $\widetilde{\Delta}$ of the cut-plane $\Delta$ and, thanks to the factor $((z_{1}\cdot z_{2})^{2}-1)^{-\frac{d-2}{4}}$, Eqs. (8.72) and (8.76) can and must be understood in this sense. In view of [19, Eq. (3.3.1.4)] one has also that
$W^{(d)}_{1-d-\lambda}(z)=W^{(d)}_{\lambda}(z)-\frac{\cos\left(\frac{1}{2}\pi(d+2\lambda)\right)\Gamma(-\lambda)\Gamma(d+\lambda-1)}{(2\pi)^{d\over 2}}(z^{2}-1)^{-\frac{d-2}{4}}P_{\frac{d-2}{2}+\lambda}^{-\frac{d-2}{2}}(z).$ (8.78)
The second term at the rhs in this relation is regular on the cut $z\in[-1,1]$ and therefore does not contribute to the commutator; this means that the two solutions are equivalent at short distances. On the other hand, since it grows the faster the larger is $|\lambda+(d-2)/2|$ (see [19, Eqs. (3.9.2))], the two-solution have drastically different long range behaviors.
The question arises whether and when they are both acceptable or not [9]. The positive definiteness condition gives the answer. We will come back on this point later.
@@
The Schwinger function (otherwise called the Euclidean propagator) is the restriction of the maximally analytic two-point function to the Euclidean Lobachevsky manifold extended to coincident points. Choosing the points in Eq. (8.72) as follows
@@ 9.2 General case: relative homology
Comparing this alternative representation of the two-point function with Eq. (9.92) we deduce a special case of a known (see [19, Eq. 45, page 136]) and useful identity which we reproduce here for the reader’s convenience:
@@ 12 The Poincaré patch $AdS_{M}$
In this section and the ones that follow, we perform a decomposition of the plane waves (7.70) and diagonalize the two-point function (1.1) in the Poincaré coordinate system. This approach paves the way for novel possibilities to compute Feynman diagrams directly in the real AdS space, as opposed to working in its Euclidean counterpart as is usually done.
We denote by $AdS_{M}$ ($M$ stands for Minkowski) the region of the $d$-dimensional manifold $AdS_{d}$ (or, what is the same, of its covering $\widetilde{}AdS_{d}$) foliated by the family hyperplanes
@@
The following system parametrizes the leaves by Lorentzian coordinates
@@
are the Lorentzian scalar product and interval in a (d1)(d-1)-dimensional Minkowski space $M_{d-1}$.
In the Poincaré coordinates the AdS scalar product and interval are written as follows:
@@
This equation tells us that the Lorentzian geometry of the leaves is coherent with that of the AdS ambient space i.e. the causal ordering in any hyperplane $\Pi_{u}$ can be understood equivalently either in the Minkowskian sense of $\Pi_{u}$ or in the sense of the ambient AdS universe. The manifold $AdS_{M}$ is strongly causal; the price to pay is that it is geodesically incomplete.
@@
This means the following: let us consider a partial complexification of $AdS_{M}$ where the Poincaré coordinates $\mathbf{x}^{\mu}$ in Eq. (12.144) are extended to complex values, while the coordinate $u$ remains real and positive:
@@
The leaves at fixed $u>0$ are copies of $M_{d-1}^{(c)}$. Suppose now that ${\rm z}={\rm x}+i{\rm y}$ and that ${\rm y}^{2}_{M}>0$, i.e. $\rm z$} belongs to either $T^{+}$ or $T^{-}$; it follows that
@@
Eq. (4.32) then implies that the Minkowskian tubes $T_{+}$ and $T_{-}$ of the leaves are contained into $\cal Z$$_{+}$ and, respectively $\cal Z$$_{-}$ (see [11] for more details). This property has also an important technical value; in particular it allows to compute the Fourier transforms of the waves w.r.t. the coordinates ${\rm x}^{\mu}$; we will do it in the next sections.
The last very important geometrical fact is that the above partial complexification already contains the whole Euclidean manifold $EAdS_{d}$.
@@
The invariant measure on $\gamma_{P}^{(\pm)}$ is $d\mu(\xi(\eta))=d\eta$ and the scalar product $x\cdot\xi$} take the form
$x(t,u)\cdot\xi(\eta)=\frac{u}{2}-\frac{\left({{\rm x}}-{\eta}\right)^{2}}{2u},\ \ \ \ \ \ \xi(\eta)\in\gamma_{P}^{(+)}$ (12.154)
@@ 13 Waves and two-point functions in Poincaré coordinates: study of the two dimensional case
The two-dimensional case is again simpler but represents a valuable training ground for solving the general case which is much harder.
In spacetime dimension $d=2$ the leaves $\Pi_{u}$ are time-like parabolae parametrized just by one time coordinate (i.e. $x^{0}=t/u$, see Eq. (12.144)). $\gamma^{+}_{P}$ is also parametrized by only one variable $\eta^{0}=\eta$ dual to the time $t$.
@@
This is yet another new integral representation of the Legendre function of the second kind $Q_{\lambda}$; this representation is naturally associated with the Poincaré coordinate system.
We now proceed to separate the variables in the above expression by taking the Fourier transform of the holomorphic plane waves (7.70) w.r.t. the time coordinate $t$:
$\int(x(t-i\epsilon,u)\cdot\xi(\eta))^{\lambda}e^{i\omega t}\,dt\,=\theta(\omega)\,\frac{(2\pi)^{\frac{3}{2}}u^{\frac{1}{2}}}{{\Gamma(-\lambda)}}\,e^{i\omega\eta}\,{\omega^{-\frac{1}{2}-\lambda}}\,J_{-\frac{1}{2}-\lambda}\left(\omega u\right),$ (13.167)
$\int(x(t+i\epsilon,u)\cdot\xi(\eta))^{\lambda}e^{i\omega t}\,dt\,=\theta(-\omega)\,\frac{(2\pi)^{\frac{3}{2}}u^{\frac{1}{2}}}{{\Gamma(-\lambda)}}\,e^{i\omega\eta}\,{|\omega|^{-\frac{1}{2}-\lambda}}\,J_{-\frac{1}{2}-\lambda}\left(|\omega|u\right),$ (13.168)
valid for for $1<\mathrm{Re}\,\lambda<0$; $\theta(\omega)$ is Heaviside’s step function.
The physically relevant feature of the above Fourier transforms is that planes waves holomorphic in the backward tube of the complex AdS manifold contain only positive frequencies w.r.t. the time coordinate of the Poincaré branes and planes waves holomorphic in the forward tube only contain negative frequencies. This support property is a consequence of the geometric inclusion of the tubes described in Eq. (12.151).
Inversion allows to rewrite the integral (13.166) as follows:
@@
$=\frac{(2\pi u^{\prime})^{1/2}}{\Gamma(\lambda+1)}\int\int_{-t-u}^{{t+u}}\left(\frac{u}{2}-\frac{\left(t-{\eta}\right)^{2}}{2u}\right)^{\lambda}e^{-i\omega(t^{\prime}+is^{\prime}-\eta)}{|\omega|^{\frac{1}{2}+\lambda}J_{\frac{1}{2}+\lambda}(|\omega|u^{\prime})\theta(-\omega)}d\omega d\eta$ (13.170)
The integral over $\eta$ contains both positive and negative frequencies (no step function is present here); the result is
@@
Putting everything together we arrive at the final result:
$W^{(2)}_{\lambda}(x_{1}(t,u),z_{2}(t^{\prime}+is^{\prime},u^{\prime}))=\frac{1}{2}{(uu^{\prime})^{1/2}}\int_{0}^{\infty}e^{-i\omega(t-t^{\prime}-is^{\prime})}J_{\frac{1}{2}+\lambda}(\omega u)J_{\frac{1}{2}+\lambda}(\omega u^{\prime})d\omega.$ (13.173)
This expression has been proven under the condition $1<\mathrm{Re}\,\lambda<0$ but is valid more generally in the sense of distributions. It provides a diagonalization of the two-point function which, for real $\lambda$, manifestly exhibits its positive-definiteness where it converges.
By comparing this expression with the one obtained by inserting the Fourier representation of both the first and the second factor in Eq. (13.166) we get as a bonus a time-honored formula of Sonine which we do not reproduce here [21].

### v2

@@ 7.1 Plane waves again
We are now in position to give a global definition of the anti-de Sitter plane waves. The same problem has been solved long ago in the de Sitter case [13, 14] (see Appendix B). Given any complex number $\lambda$ the natural domains of definition for plane waves in the AdS universe are ${\cal Z}_{+}\times\overline{{\cal C}}_{-}$ and ${\cal Z}_{-}\times\overline{{\cal C}}_{+}$:
@@
A previous attempt [32] to introduce global waves in AdS was based on a local $i\epsilon$-prescription. Here, in their global domains of definition, the waves are univalued, holomorphic solutions of the Klein-Gordon equation with complex squared mass $m^{2}_{\lambda}=\lambda(\lambda+d-1).$
@@ 8 Klein-Gordon fields
Klein–Gordon fields provide the simplest example of the analytic structure described in section 5. Their two-point functions are distributions on $\widetilde{AdS_{d}}\times\widetilde{AdS_{d}}$ satisfying the AdS Klein–Gordon equation with respect to both variables:
$\left(\square_{x_{1}}+m_{\lambda}^{2}\right)\mathcal{W}(x_{1},x_{2})=\left(\square_{x_{2}}+m_{\lambda}^{2}\right)\mathcal{W}(x_{1},x_{2})=0,$ (8.76)
where the (complex) squared mass is parametrized by the complex number $\lambda$ as in eq. (6.51).
For any such $\lambda$ there exist two invariant maximally analytic solutions satisfying all the properties listed in section 5, except possibly positive definiteness (unitarity). They are function of the complex variable $z=z_{1}\cdot z_{2}$ and are mapped into each other by the involution (6.54):
$W^{(d)}_{\lambda}(z_{1},z_{2})$ $=W^{(d)}_{\lambda}(z)=\frac{e^{-i\pi\frac{d-2}{2}}}{(2\pi)^{\frac{d}{2}}}\big(z^{2}-1\big)^{-\frac{d-2}{4}}Q^{\,\frac{d-2}{2}}_{\frac{d-2}{2}+\lambda}(z)$ (8.77)
$=\frac{z^{1-d-\lambda}\Gamma(d-1+\lambda)\,{}_{2}F_{1}\!\left(\frac{d-1+\lambda}{2},\frac{d+\lambda}{2};\frac{d+1+2\lambda}{2};z^{-2}\right)}{2^{d+\lambda}\pi^{(d-1)/2}\Gamma\!\left(\frac{d+1}{2}+\lambda\right)}$ (8.78)
$W^{(d)}_{1-d-\lambda}(z_{1},z_{2})$ $=\frac{e^{-i\pi\frac{d-2}{2}}}{(2\pi)^{\frac{d}{2}}}\big(z^{2}-1\big)^{-\frac{d-2}{4}}Q^{\,\frac{d-2}{2}}_{-\frac{d}{2}-\lambda}(z).$ (8.79)
By using the parameter $\nu$ introduced in eq. (6.56), the above solutions can also be written as
$W^{(d)}_{\nu(\lambda)}(z_{1},z_{2})$ $=\frac{e^{-i\pi\frac{d-2}{2}}}{(2\pi)^{\frac{d}{2}}}\big(z^{2}-1\big)^{-\frac{d-2}{4}}Q^{\,\frac{d-2}{2}}_{-\frac{1}{2}+\nu}(z)$ (8.80)
$=\frac{\Gamma\!\left(\frac{d-1}{2}+\nu\right)(2z)^{-\frac{d-1}{2}-\nu}\,{}_{2}F_{1}\!\left(\frac{d-1}{4}+\frac{\nu}{2},\frac{d+1}{4}+\frac{\nu}{2};\nu+1;\frac{1}{z^{2}}\right)}{2\pi^{\frac{d-1}{2}}\Gamma(\nu+1)},$ (8.81)
$W^{(d)}_{-\nu(\lambda)}(z_{1},z_{2})$ $=\frac{e^{-i\pi\frac{d-2}{2}}}{(2\pi)^{\frac{d}{2}}}\big(z^{2}-1\big)^{-\frac{d-2}{4}}Q^{\,\frac{d-2}{2}}_{-\frac{1}{2}-\nu}(z).$ (8.82)
As expected, the involution acts simply by changing the sign of the parameter $\nu$.
@@
In the above formulae, and in what follows, $P^{\alpha}_{\beta}$ and $Q^{\alpha}_{\beta}$ denote the associated Legendre functions, which are meromorphic in their indices [18]. As functions of the complex variable $z$, they are analytic in the cut plane
$\Delta_{2}=\mathbf{C}\setminus[-\infty,1].$ (8.83)
As explained in Section 4, the AdS scalar product $z_{1}\!\cdot\!z_{2}$ lifts to a map
$\widetilde{\mathcal{Z}}_{-}\times\widetilde{\mathcal{Z}}_{+}\to\widetilde{\Delta}_{1},$
(the covering of the cut plane $\Delta_{1}=\mathbb{C}\setminus[-1,1]).$ Equations (8.77)–(8.82) are to be understood in this sense.
The Legendre function of the second kind $Q^{\,\frac{d-2}{2}}_{-\frac{1}{2}+\nu}(z)$ (or equivalently $Q^{-\frac{d-2}{2}}_{-\frac{1}{2}+\nu}(z)$), multiplied by the factor $(z-1)^{-\frac{d-2}{4}}(z+1)^{-\frac{d-2}{4}}$, are holomorphic in $\widetilde{\Delta}_{1}$, a domain naturally adapted to the anti-de Sitter geometry. Points on the cut [1,1][-1,1] of the fundamental sheet $\Delta_{1}$ correspond to pairs of timelike separated events on $AdS_{d}$}.
In contrast, the following Legendre functions of the first kind [^3] In this case the sign of the upper index is uniquely fixed; these functions are also known as Gegenbauer functions [18]., multiplied by the same prefactor
$(z-1)^{-\frac{d-2}{4}}(z+1)^{-\frac{d-2}{4}}P^{-\frac{d-2}{2}}_{-\frac{1}{2}+i\kappa}(z),\qquad\kappa\in\mathbb{C},$ (8.84)
possess analyticity properties that are naturally adapted to the Lorentzian geometry of the de Sitter manifold [13, 14] (see Appendix B) as they are analytic in the cut plane
$\Delta=\mathbf{C}\setminus(-\infty,-1].$ (8.85)
The prefactor exactly compensates the singularity of the Legendre function and the product i regular there. Points lying on the cut $\infty,-1]$ correspond to pairs of timelike separated real events on the de Sitter manifold $dS_{d}$.
The symmetry property of the Legendre functions of the first kind,
$P^{\mu}_{-\frac{1}{2}-i\kappa}(z)=P^{\mu}_{-\frac{1}{2}+i\kappa}(z),\qquad\kappa\in\mathbb{C},\;z\in\Delta.$ (8.86)
does not hold for the Legendre functions of the second kind. As a consequence, in the AdS case two distinct two-point functions can arise for a given mass, although they are not independent. Indeed, using [18, Eq. (3.3.1.4)], one obtains the identity
$Q_{-\frac{1}{2}-\nu}^{\mu}(z)=Q_{-\frac{1}{2}+\nu}^{\mu}(z)+e^{i\pi\mu}\sin(\pi\nu)\Gamma\!\left(\tfrac{1}{2}+\mu-\nu\right)\Gamma\!\left(\tfrac{1}{2}+\mu+\nu\right)P_{-\frac{1}{2}+\nu}^{-\mu}(z).$ (8.87)
This relation implies that
$W^{(d)}_{-\nu}(z)=W^{(d)}_{\nu}(z)+\frac{\sin(\pi\nu)\Gamma\left(\frac{d-1}{2}+\nu\right)\Gamma\left(\frac{d-1}{2}-\nu\right)}{(2\pi)^{\frac{d}{2}}}(z^{2}-1)^{-\frac{d-2}{4}}P_{-\frac{1}{2}+\nu}^{-\frac{d-2}{2}}(z).$ (8.88)
The second term on the right-hand side is regular on the AdS cut $z\in[-1,1]$ and therefore does not contribute to the commutator. Thus the two solutions coincide at short distances. However, since this term grows increasingly rapidly as $|\nu|$ becomes large (see [18, Eqs. (3.9.2)]), the two solutions exhibit markedly different long-distance behavior.
The question of whether both solutions are physically admissible has been solved long ago: the answer is provided by the requirement of positive definiteness (unitarity) [9]. An alternative viewpoint invoke the possible self-adjoint extensions of a certain Schrödinger operator [25, 26].
@@
The Schwinger function is the restriction of the maximally analytic two-point function to non-coincident Euclidean points:
$S(z_{1}^{\scriptscriptstyle{E}},{z_{2}^{\scriptscriptstyle{E}})=\left.W(z_{1},z_{2})\right|_{EAdS_{d}\times EAdS_{d}\setminus\delta}}$ (8.89)
where $\delta$ is the diagonal. The Euclidean propagator coincides with the two-point Schwinger function away from coincident points and is its distributional extension to the diagonal. Choosing the points in Eq. (8.77) as follows
@@ 9.2 General case: relative homology
Comparing this alternative representation of the two-point function with Eq. (9.103) we deduce a special case of a known (see [18, Eq. 45, page 136]) and useful identity which we reproduce here for the reader’s convenience:
@@ 12 The Poincaré patch $AdS_{M}$
In this and the subsequent sections, we decompose the plane waves (7.75) and diagonalize the two-point function (1.1) using the Poincaré coordinate system. This framework opens up new opportunities to compute Feynman diagrams directly in Lorentzian AdS space, rather than relying on the conventional Euclidean approach.
We denote by $AdS_{M}$ (with $M$ standing for Minkowski) the region of the $d$-dimensional manifold $AdS_{d}$} - or equivalently, of its universal covering space $\widetilde{AdS_{d}} -$ that is foliated by the family of hyperplanes
@@
The following system parametrizes the leaves by Lorentzian (Poincaré) inertial coordinates
@@
are the scalar product and the interval in a (d1)(d-1)-dimensional Minkowski space $M_{d-1}$. In Poincaré coordinates, the AdS scalar product and interval are expressed as follows:
@@
This equation shows that the Lorentzian geometry of the leaves is consistent with that of the ambient AdS space; that is, the causal ordering within any hyperplane $\Pi_{u}$ can be understood either in the Minkowskian sense of $\Pi_{u}$ or in the context of the ambient AdS spacetime. The manifold $AdS_{M}$ is strongly causal, though this comes at the cost of geodesic incompleteness.
@@
This can be stated as follows: consider a partial complexification of $AdS_{M}$ in which the Poincaré coordinates $\mathbf{x}^{\mu}$ in Eq. (12.155) are extended to complex values, while the coordinate $u$ remains real and positive:
@@
The leaves at fixed $u>0$ are copies of $M_{d-1}^{(c)}$. Suppose now that ${\rm z}={\rm x}+i{\rm y}$ and that ${\rm y}^{2}_{M}>0$, i.e. $\rm z$} belongs to either $T^{+}$ or $T^{-}$; it follows immediately that
@@
equation (4.33) then implies that the Minkowskian tubes $T_{+}$ and $T_{-}$ of the leaves are contained within $\mathcal{Z}_{+}$ and, respectively, $\mathcal{Z}_{-}$ (see [10] for further details). This property also has important technical significance: in particular, it allows the computation of Fourier transforms of the waves with respect to the coordinates $\mathbf{x}^{\mu}$, which we will carry out in the following sections.
Finally, a crucial geometrical observation is that the above partial complexification already encompasses the entire Euclidean manifold $EAdS_{d}$.
@@
The invariant measure on $\gamma_{P}^{(\pm)}$ is $d\mu(\xi(\eta))=d\eta$ and the scalar product $x\cdot\xi$} takes the forms
$x(t,u)\cdot\xi(\eta)=\frac{u}{2}-\frac{\left({{\rm x}}-{\eta}\right)^{2}}{2u},\ \ \ \ \ \ \xi(\eta)\in\gamma_{P}^{(+)},$ (12.165)
@@ 13 Waves and two-point functions in Poincaré coordinates: study of the two dimensional case
The two-dimensional case is simpler but still provides a valuable training ground for tackling the more general and challenging higher-dimensional case.
In spacetime dimension $d=2$, the leaves $\Pi_{u}$ are time-like parabolae, each parametrized by a single time coordinate ${\rm x}^{0}=t$, i.e.,
$x^{0}=\frac{t}{u}\,$, (13.175)
(see Eq. 12.155). Similarly, $\gamma^{+}_{P}$ is parametrized by a single dual variable $\eta^{0}=\eta$, which is conjugate to the time $t$.
@@
This provides yet another integral representation of the Legendre function of the second kind, $Q_{\lambda}$, which is naturally associated with the Poincaré coordinate system.
We now proceed to separate variables in the above expression by taking the Fourier transform of the holomorphic plane waves (7.75) with respect to the time coordinate $t$. The result is
$\int(x(t-i\epsilon,u)\cdot\xi(\eta))^{\lambda}e^{i\omega t}\,dt\,=\,\theta(\omega)\,\frac{(2\pi)^{\frac{3}{2}}u^{\frac{1}{2}}}{\Gamma(-\lambda)}\,e^{i\omega\eta}\,\omega^{-\frac{1}{2}-\lambda}\,J_{-\frac{1}{2}-\lambda}(\omega u),$ (13.179)
$\int(x(t+i\epsilon,u)\cdot\xi(\eta))^{\lambda}e^{i\omega t}\,dt\,=\,\theta(-\omega)\,\frac{(2\pi)^{\frac{3}{2}}u^{\frac{1}{2}}}{\Gamma(-\lambda)}\,e^{i\omega\eta}\,|\omega|^{-\frac{1}{2}-\lambda}\,J_{-\frac{1}{2}-\lambda}(|\omega|u),$ (13.180)
valid for $1<\mathrm{Re}\,\lambda<0$, where $\theta(\omega)$ is Heaviside’s step function.
A physically relevant feature of the above Fourier transforms is that plane waves holomorphic in the backward tube of the complex AdS manifold contain only positive frequencies with respect to the time coordinate of the Poincaré branes, whereas plane waves holomorphic in the forward tube contain only negative frequencies. This support property follows from the geometric inclusion of the tubes described in Eq. (12.162).
Inversion then allows us to rewrite the integral (13.178) as follows:
@@
$=\frac{(2\pi u^{\prime})^{1/2}}{\Gamma(\lambda+1)}\int\int_{-t-u}^{{t+u}}\left(\frac{u}{2}-\frac{\left(t-{\eta}\right)^{2}}{2u}\right)^{\lambda}e^{-i\omega(t^{\prime}+is^{\prime}-\eta)}{|\omega|^{\frac{1}{2}+\lambda}J_{\frac{1}{2}+\lambda}(|\omega|u^{\prime})\theta(-\omega)}d\omega d\eta.$ (13.182)
The integral over $\eta$ contains both positive and negative frequencies (no step function is present here):
@@
Putting everything together, we arrive at the final result in the form of a Hankel representation:
$W^{(2)}_{\lambda}\big(x_{1}(t,u),z_{2}(t^{\prime}+is^{\prime},u^{\prime})\big)$ $=$ $\frac{1}{2}\,(uu^{\prime})^{1/2}\int_{0}^{\infty}e^{-i\omega(t-t^{\prime}-is^{\prime})}\,J_{\frac{1}{2}+\lambda}(\omega u)\,J_{\frac{1}{2}+\lambda}(\omega u^{\prime})\,d\omega,$ (13.185)
which has been proven under the condition $1<\mathrm{Re}\,\lambda<0$, but is valid more generally in the sense of distributions. It provides a diagonalization of the two-point function which, for real $\lambda$, manifestly exhibits its positive-definiteness in the region where it converges.
By comparing this expression with the one obtained by inserting the Fourier representation of both the first and second factors in Eq. (13.178), we also recover a classical formula of Sonine, which we do not reproduce here [19].